# Bayesian mixture modeling using a mixture of finite mixtures with normalized inverse Gaussian weights

### v1

@@ 4.1.2 Thyroid Data
We apply the proposed method to famous thyroid data. The data is available from the R package mclust and is well known as benchmark data for clustering. The sample size is $215$ and the dimension of the data is $6$. The thyroid disease of each patient is included and classified into three categories: Normal, hypo and hyper. The number of diseases is $150$, $30$ and $35$, respectively. Using these labels as true labels, the main interest is the accuracy of the clustering and whether the number of components is estimated to be $3$.
We use the same model and shape parameters as in Section 4.1.1. We independently run the $5$ MCMC chain with different initial values. The number of iterations for each chain is $20000$ and finally we get the $10000$ samples. We compare MFM-Inv-Ga and MFM-Ga through the posterior mean of $M$, the posterior probability $\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$ and the posterior mean of the rand index (RI). Table 4 shows that the result of the real data analysis is similar to that of Section 4.1.1. MFM-Inv-Ga provides a reasonable estimate of $M$ for all shape parameters by not producing empty components. However, MFM-Ga overestimates $M$. Furthermore, MFM-Inv-Ga is slightly more accurate than MFM-Ga.
Table 4: Posterior means of $M$ ($\hat{M}$), posterior probabilities of $M_{na}=0$ ($\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$) and posterior means of rand index (RI) for thyroid data.
MFM-Inv-Ga
$\hat{M}$ $\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$ RI
$\alpha=1$ $3.200$ $1.000$ $0.895$
$\alpha=0.2$ $3.213$ $0.994$ $0.895$
$\alpha=10^{-1}$ $3.225$ $0.987$ $0.894$
$\alpha=10^{-2}$ $3.222$ $0.987$ $0.895$
$\alpha=10^{-3}$ $3.489$ $0.960$ $0.893$
MFM-Ga
$\hat{M}$ $\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$ RI
$\gamma=1$ $3.670$ $0.962$ $0.893$
$\gamma=0.2$ $4.944$ $0.434$ $0.890$
$\gamma=10^{-1}$ $4.945$ $0.394$ $0.893$
$\gamma=10^{-2}$ $7.425$ $0.067$ $0.893$
$\gamma=10^{-3}$ $7.626$ $0.046$ $0.895$
@@ 4.2 Density estimation
As seen in the previous subsection, the main difference between the MFM-Inv-Ga and MFM-Ga methods is the frequency of occurrence of the empty components. The MFM-Ga method can achieve very high clustering accuracy by setting a small $\gamma$ and allowing empty components. However, small $\gamma$ not only makes the model difficult to interpret but also degrades the predictive accuracy of the model. In this subsection, we examine this phenomenon through density estimation using predictive distributions.
@@
We used the famous galaxy dataset, which is a small data set consisting of $82$ velocities (km/sec) of different galaxies. The data is widely used in nonparametric Bayesian statistics as a benchmark for density estimation and cluster analysis. The details of the data are found in Roeder (1990). We set $M-1|\Lambda\sim\text{Poisson}(\Lambda)$ and $\Lambda\sim\text{Ga}(1,1/5)$. We also employ the univariate normal kernel $N(y|\mu,\sigma^{2})$, and the normal–inverse gamma conjugate prior $N(\mu|m_{0},\sigma^{2}/\tau)\times\text{IG}(\sigma^{2}|c_{0},C_{0})$ as the prior of the parameters in the kernel. Moreover, we assume that the prior of $C_{0}$ is $\text{Gamma}(d_{0},D_{0})$, and we set $m_{0}=(\max(\mathrm{data})+\min(\mathrm{data}))/2,\ d_{0}=0.2,\ D_{0}=10/(\max(\mathrm{data})+\min(\mathrm{data}))^{2}$ as Richardson and Green (1997). The parameter $\tau$ controls the smoothness of the estimated density function. We assume that the prior of the smoothing parameter $\tau$ is $\text{Gamma}(w,W)$, where $w=0.5$ and $W=50$ as Escobar and West (1995). Parameters $\tau$ and $C_{0}$ should be carefully learned from the data, since density estimation is sensitive to their choice. The MCMC iterations are $100000$ and the first half of the $90000$ samples are not used as a burn-in period. The shape parameters are the same as in Section 4.1. We evaluated the result of density estimation and posterior probabilities of $k$, $M$, and $M_{na}$ for each shape parameter.
Figure 3: Results of density estimation for galaxy data using MFM-Inv-Ga (left) and MFM-Ga (right).
@@
We show the estimated density functions using the posterior means in Figure 3. From the figure, it is observed that the shapes of the estimated densities using the MFM-Inv-Ga do not depend on the choice of the shape parameter $\alpha$. However, results using the MFM-Ga method seem to be strongly influenced by the choice of the shape parameter $\gamma$. For $\gamma=10^{-2}$ and $10^{-3}$, MFM-Ga cannot capture two large peaks in the middle.
Table 5: Posterior probabilities of the number of components $M$ and clusters $k$ for Galaxy data.
MFM-Inv-Ga
$M\leq 3$ $M=4$ $M=5$ $M=6$ $M=7$ $M=8$ $M=9$ $M\geq 10$
$\alpha=1.0$ $0.000$ $0.089$ $0.224$ $0.31$ $0.170$ $0.097$ $0.054$ $0.056$
$\alpha=0.2$ $0.136$ $0.194$ $0.171$ $0.152$ $0.114$ $0.086$ $0.054$ $0.091$
$\alpha=10^{-1}$ $0.071$ $0.096$ $0.154$ $0.144$ $0.137$ $0.102$ $0.079$ $0.216$
$\alpha=10^{-2}$ $0.039$ $0.078$ $0.119$ $0.151$ $0.123$ $0.102$ $0.081$ $0.308$
$\alpha=10^{-3}$ $0.057$ $0.111$ $0.170$ $0.153$ $0.122$ $0.101$ $0.071$ $0.217$
@@
The number of clusters in the galaxy data has been reported as $5$ or $6$ in existing studies. From table 5, in MFM-Inv-Ga, the posterior distributions of $k$ have large probabilities at $k=5$ and $6$ for all $\alpha$ and each posterior distribution induced by MFM-Inv-Ga is more similar than by MFM-Ga. For $M$, both MFM-Inv-Ga and MFM-Ga overestimate when $\alpha$ and $\gamma$ are small. This can be seen in Table 6. The reason is why the data size $n=82$ is small and $M_{na}$ is easily produced. Tables 5 and 6 show that the degree of overestimation is much smaller for MFM-Inv-Ga than for MFM-Ga.
Table 6: Posterior probabilities of the number of empty components is equal to zero for Galaxy data
$\mathbb{P}(M_{na}=0|y_{1},\dots,y_{n})$
$\alpha$ and $\gamma$ $1$ $0.2$ $10^{-1}$ $10^{-2}$ $10^{-3}$
MFM-Inv-Ga $0.854$ $0.654$ $0.480$ $0.347$ $0.443$
MFM-Ga $0.648$ $0.081$ $0.040$ $0.010$ $0.002$
Focusing on the number of clusters $k$, it is interesting that MFM-Inv-Ga is more robust for shape parameter estimates $k$ than MFM-Ga. This suggests that the prior distribution of $k$ based on MFM-Inv-Ga is less informative than MFM-Ga, i.e., the same relationship holds for MFM-Inv-Ga and MFM-Ga as for normalized inverse Gaussian process and Dirichlet process. In Lijoi et al. (2005), the prior of $k$ induced by a normalized inverse Gaussian process is not more informative for the precision parameter than the Dirichlet process. However, the prior $k$ for the normalized inverse Gaussian process can be written in closed form, while that for MFM-Inv-Ga is given in complex integral form.
The shape parameter of MFM-Ga should be chosen carefully because it has a significant impact on clustering, density estimation, and the appearance of empty components. However, MFM-Inv-Ga is much more robust than MFM-Ga with respect to the choice of the shape parameter. Hence, MFM-Inv-Ga is superior to MFM-Ga in that it is much easier to use than MFM-Ga.

### v2

@@ 4.1.2 Thyroid data
We apply the proposed method MFM-IGau and DMFM-IGau to famous thyroid data. The main purpose of this analysis is to infer the number of cluster $k$ and the number of components $M$ and examine the impact of choice of $q_{M}$. The thyroid data are available from the R package mclust and is well known as benchmark data for clustering. The sample size is $215$ and the dimension of the data is $6$. The thyroid disease of each patient is included and classified into three categories: Normal, hypo and hyper. The number of diseases are $150$, $30$ and $35$, respectively. Using these labels as true labels, the main interest is whether the number of components is estimated to be $3$. We use priors Poisson distribution $\mathrm{Poisson}(1)$, geometric distribution $\mathrm{Geo}(0.1)$ and beta negative binomial distribution $\mathrm{BNB}(1,4,3)$ for $M$, and $F$-distribution $F(6,3)$ for $\alpha$.
Table 1 shows that the posterior modes for MFM-IGau and DMFM-IGau are all equal to $3$. These results are consistent with those obtained from DMFM-Ga. In addition, the table indicates that both the first and third quartiles are $3$, and that the posterior probabilities of $M_{na}=0$ are very close to 1. This implies that the difference between the posterior number of components and the posterior number of clusters obtained by MFM-IGau and DMFM-IGau is extremely small. Furthermore, the interquartile range of $M$ is tighter than that reported in fruhwirth2021generalized, providing stronger evidence that $M=3$ under our proposed methods than under DMFM-Ga. In addition, the overall results do not depend on the choice of the prior $q_{M}$, as is the case for DMFM-Ga.
Table 1: Posterior summaries of $k$ and $M$ for the thyroid data. The leftmost values represent the posterior modes, the values in brackets denote the first and third quartiles, and the rightmost values correspond to the posterior probabilities of $M_{na}=0$.
$p(k\mid\mathrm{data})$
$\mathrm{Poi}(1)$ $\mathrm{Geo}(0.1)$ $\mathrm{BNB}(1,4,3)$
MFM-IGau 33 $[3,3]$ 33 $[3,3]$ 33 $[3,3]$
DMFM-IGau 33 $[3,3]$ 33 $[3,3]$ 33 $[3,3]$
$p(M\mid\mathrm{data})$
$\mathrm{Poi}(1)$ $\mathrm{Geo}(0.1)$ $\mathrm{BNB}(1,4,3)$
MFM-IGau 33 $[3,3]$ $0.997$ 33 $[3,3]$ $0.980$ 33 $[3,3]$ $1.000$
DMFM-IGau 33 $[3,3]$ $0.993$ 33 $[3,3]$ $0.973$ 33 $[3,3]$ $0.986$
@@ 4.2 Density estimation
In this section, we compare MFM-IGau and MFM-Ga using the Galaxy data. The aim of this analysis is to examine how sensitive the posterior distribution of $k$ and the density estimation based on the predictive distribution are to the choice of shape parameters.
We used the famous galaxy dataset, which is a small data set consisting of $82$ velocities (km/sec) of different galaxies. The data are widely used in nonparametric Bayesian statistics as a benchmark for density estimation and cluster analysis. The details and a more comprehensive analysis of the data were given in roeder1990density and grun2022many.
@@
We set $M-1\mid\Lambda\sim\text{Poisson}(\Lambda)$ and $\Lambda\sim\text{Ga}(1,1/5)$. We also employ the univariate normal kernel $N(y\mid\mu,\sigma^{2})$, and the normal-nverse gamma conjugate prior in Section 3.4.1: $N(\mu\mid m_{0},\sigma^{2}/\tau)\times\text{IG}(\sigma^{2}\mid c_{0},C_{0})$ as the prior of the parameters in the kernel. Moreover, we assume that the prior of $C_{0}$ is $\text{Gamma}(d_{0},D_{0})$, and we set $m_{0}=(\max(\mathrm{data})+\min(\mathrm{data}))/2,\ d_{0}=0.2,\ D_{0}=10/(\max(\mathrm{data})+\min(\mathrm{data}))^{2}$ as richardson1997bayesian. The parameter $\tau$ controls the smoothness of the estimated density function. We assume that the prior of the smoothing parameter $\tau$ is $\text{Gamma}(w,W)$, where $w=0.5$ and $W=50$ as escobar1995bayesian. Parameters $\tau$ and $C_{0}$ should be carefully learned from the data, since density estimation is sensitive to their choice. MCMC was run for 100,000 iterations with a burn-in of 90,000. Each value of $\alpha=1.0,0.2,10^{-1},10^{-2},10^{-3}$ is considered, and $\gamma$ is set to $\gamma=\alpha$ in each case. Note that the first and second moments of $\mathrm{IGau}(\alpha,1)$ and $\mathrm{Gamma}(\gamma,1)$ are both equal to $\alpha$ and $\gamma$, respectively; thus, this corresponds to a moment-matching setting. We evaluated the result of density estimation and posterior probabilities of $k$ for each shapes.
@@
We show the estimated density functions using the posterior means in Figure 3. From the figure, it is observed that the shapes of the estimated densities using the MFM-IGau do not depend on the choice of the shape parameter $\alpha$. However, results using the MFM-Ga method seem to be strongly influenced by the choice of the shape parameter $\gamma$. For $\gamma=10^{-2}$ and $10^{-3}$, MFM-Ga cannot capture two large peaks in the middle.
Figure 3: Results of density estimation for Galaxy data using MFM-IGau (left) and MFM-Ga (right).
The number of clusters in the galaxy data has been reported as $5$ or $6$ in existing studies. From Table 2, in MFM-IGau, the posterior distributions of $k$ have large probabilities at $k=5$ and $6$ for all $\alpha$, and the posterior distributions of $k$ under MFM-IGau are more similar to each other across different value of $\alpha$ than those under MFM-Ga. Focusing on the number of clusters $k$, it is interesting that MFM-IGau is more robust than MFM-Ga with respect to the choice of the shape parameter in estimating $k$. This suggests that the prior distribution of $k$ based on MFM-IGau is less informative than MFM-Ga, i.e., the same relationship holds for MFM-IGau and MFM-Ga as for normalized inverse Gaussian process and Dirichlet process.
Table 2: Posterior probabilities of the number of clusters $k$ for Galaxy data.
MFM-IGau
@@
The shape parameter of MFM-Ga should be chosen carefully because it has a significant impact on clustering, density estimation, and the appearance of empty components. However, MFM-IGau is much more robust than MFM-Ga with respect to the choice of the shape parameter. This is consistent with the earlier remark that the normalized inverse Gaussian distribution is less informative and more robust to parameter specification than the Dirichlet distribution, leading to more stable inference. Note that results do not differ substantially when information measures such as the expected Gini index are matched.